OpenStudy (anonymous):

how do you solve cos(2x)*2=0 have to do rolle's theorem all assumptions satisfied and this the derivative of the functionhow do you solve cos(2x)*2=0 have to do rolle's theorem all assumptions satisfied and this the derivative of the function@Mathematics

OpenStudy (anonymous):

sin(2x) + C has the same value for x = 0 and x = pi, also for x = pi and x = 2pi, etc. so therefore the derivative of sin(2x) + C (which is 2*cos(2x)) must be 0 between those intervals, so it must be 0 between [0, pi], between [pi, 2pi] etc.

OpenStudy (phi):

If you are asking what is x for 2cos(2x)=0 , then it is cos(2x)= 0 (divide by 2) 2x= acos(0) (take the inverse cosine) 2x= pi/2 (principal value, there are lots of solutions) x= pi/4